Go to the page in the study guide that shows the entire glycolysis pathway. If you haven't already done so, go through and indicate every reaction that consumes an ATP (ATP → ADP) and every reaction that creates an ATP (ADP → ATP). Look at the entire pathway (ie. from glucose to 2 molecules of pyruvate) and fill in the correct numbers for:

How many total ATP are consumed:   

How many total ATP are produced:   

How many net ATP are produced:
Answer #1

A.preparatory phase

First priming reaction

Glucose to glucose 6 phosphate

one ATP used

Second priming reaction fructose 6 phosphate is converted to fructose 1,6 biphosphate

One ATP used

B. Pay-off phase

1,3 biphosphoglycerate is concerned to 3 phospho glycerate

First ATP forming reaction

Substrate level phosphorylation

Two molecules of ATP formed

Then phospho enol.pyruvate is converted to pyruvate

Substrate level phosphorylation

Two molecules of ATP formed

Total ATP production = 4 ATP

Net yield of ATP= 4-2= 2 ATP

In addition 2 NADH MOLECULES are formed

= 2*3= 6 ATP

Net energy yield = 2+6= 8 ATP